Suppose that a ball is dropped from a tower 70 m high. How far will it fall after 3 seconds?
1
Expert's answer
2021-09-14T09:35:19-0400

After 3 seconds, the ball will be


"h=\frac{gt^2}{2}=\frac{9.8·3^2}2=44.1\text{ m}"

below the point where it was dropped.